The \underline{MIT Lindy Hop Society} hosts

Swing Dancing

We have not yet decided on event dates and format.*

All lessons and dances will be free. Beginners welcome, no partner necessary.

For more information, or to be subscribed to our mailing list, send mail to swing@mit.edu or see

our web page at \underline{http://mit.edu/swing}

*Before the pandemic, we held a weekly lesson and dance, every Wednesday evening.

Swing Dancing

We have not yet decided on event dates and format.*

All lessons and dances will be free. Beginners welcome, no partner necessary.

For more information, or to be subscribed to our mailing list, send mail to swing@mit.edu or see

our web page at \underline{http://mit.edu/swing}

*Before the pandemic, we held a weekly lesson and dance, every Wednesday evening.

Swing Dancing

We have not yet decided on event dates and format.*

All lessons and dances will be free. Beginners welcome, no partner necessary.

For more information, or to be subscribed to our mailing list, send mail to swing@mit.edu or see

our web page at \underline{http://mit.edu/swing}

*Before the pandemic, we held a weekly lesson and dance, every Wednesday evening.

Swing Dancing

We have not yet decided on event dates and format.*

All lessons and dances will be free. Beginners welcome, no partner necessary.

For more information, or to be subscribed to our mailing list, send mail to swing@mit.edu or see

our web page at \underline{http://mit.edu/swing}

*Before the pandemic, we held a weekly lesson and dance, every Wednesday evening.